please help,

solve the following:
7|z+6|+8 is less than or equal to 57

Answer by stanbon(75887)

7|z+6|+8 <= 57
Subtract 8 from both sides:
7|z+6| <= 49
Divide both sides by 7:
|z+6| <= 7
Rewrite as:
-7 <= z+6 <= 7
Subtract 6 along the line to get:
-13 <= z <= 1
===========
Cheers,
Stan H.